# Physics-Enhanced Deep Surrogate for the Phonon Boltzmann Transport Equation

### v1

@@ 2.1 Uncertainty Quantification with Deep Ensembles and Heteroskedastic Variance
Uncertainty Quantification (UQ) is essential in surrogate modeling for scientific applications [psaros2023uncertainty], especially when high-fidelity simulations are computationally expensive and data is limited [abdar2021review]. We adopt a heteroskedastic Gaussian surrogate model, where the mean $\kappa_{\theta}(x)\in\mathbb{R}^{N}$ is predicted by an ensemble of PEDS [lakshminarayanan2017simple] and the input-dependent log-variance $(\log\sigma^{2})_{\theta}(x)$ is predicted by a vanilla neural network. The model is trained using a Gaussian negative log-likelihood loss [bishop2006pattern, lakshminarayanan2017simple]: $\mathcal{L}_{\text{NLL}}=\sum_{i}\beta\cdot(\log\sigma^{2})_{\theta}(x_{i})+(1-\beta)\cdot\frac{(\kappa(x_{i})-\kappa_{\theta}(x_{i}))^{2}}{\exp((\log\sigma^{2})_{\theta}(x_{i}))}$, where the parameter $\beta\in[0,1]$ controls the trade-off between penalizing overconfident predictions and fitting the data. We train an ensemble of $M$ independently initialized surrogate models to further capture uncertainty arising from model variability, especially from the optimization of their parametric parts. The ensemble prediction is given by the empirical mean $\hat{\kappa}(x)=\frac{1}{M}\sum_{i=1}^{M}\kappa_{i}(x)$, and the total predictive variance combines the predicted variance and the variance of the predictions as $\hat{\sigma}^{2}(x)=\frac{1}{M}\sum_{i=1}^{M}\sigma_{i}^{2}(x)+\frac{1}{M}\sum_{i=1}^{M}(\kappa_{i}(x)-\hat{\kappa}(x))^{2}$. This approach robustly predicts epistemic uncertainty, supports active learning by prioritizing uncertain samples, and facilitates stochastic optimization by taking into account the confidence in the surrogate’s output rather than relying solely on point estimates.
@@ 2.2 Active Learning
PEDS has previously shown great data savings and better performance on challenging designs when paired with an Active Learning framework [Pestourie_Mroueh_Rackauckas_Das_Johnson]. We implemented an uncertainty-driven active learning pipeline using our uncertainty measure following Ref. [pestourie2020active]. We initialize a dataset with N training points, we train our surrogate and then propose M new geometries, compute their uncertainty and only compute and add to the dataset the $K\ll M$ most uncertain points. This is in contrast with other sample proposal strategies that are based upon the diversity of the samples or a mixture of diversity and uncertainty [settles2012active, shui2020deep]. In our case the measure of uncertainty is the variance as computed above. For active learning, unlike complex data assimilation problems requiring detailed distribution estimates, a coarse approximation of uncertainty is sufficient [abdar2021review].
@@ 3 Design of Porous-structure with desired effective Thermal Conductivity
At the nano-scale, heat conduction diverges from the classical Fourier Law [Chen2005, Chen2021] and assumes the form of a ballistic process rather than a diffusive one. This is because the mean free path (MFP) of heat-carrying quasi-particles in semiconductors, i.e. phonons, is comparable to the characteristic dimensions of the material. This phenomenon, often referred to as phonon-size effect [Chen2005, Tang2010], leads, in porous material, to a suppressed effective thermal conductivity compared to the one obtained with a standard Fourier solver [Song2004, Hochbaum2008, Lee2015] (Fig. 2). These effects hold great promise across multiple engineering domains, shifting the focus from bulk materials to nanostructures and their inverse design.
In recent work, direct topology‐optimization of porous nanostructures has been attempted using the phonon BTE. For example, Ref. [Wei2020] employed a genetic‐algorithm search with a gray (single‐MFP) BTE model for nanoporous graphene and found disordered pore patterns that increase thermal conductivity. However, gray (single‐relaxation‐time) models neglect the full phonon spectrum and generally overestimate conductivity. In fact, gray-model solutions can qualitatively differ from full-spectrum solutions: for instance, Ref. [iheduru2023comparison] shows that a gray BTE systematically predicts a higher effective conductivity than a mode-resolved BTE for the same nanofilm. Thus, designs based on single‐MFP models may yield very different results than those from a full phonon-spectrum BTE. The development of differentiable BTE frameworks is an active research area. The first implementation of such a framework in the context of inverse design (Ref. [romano2022inverse]) focused on the single-MFP model; only recently, a mode-resolved differentiable BTE solver (based on the relaxation-time approximation) was developed [shang2025jax]; however, this framework did not include the interpolation model needed for inverse design (Ref. [romano2022inverse]).
@@ 3.1 Problem Statement
We consider a 2D domain $\Omega=[-L/2,L/2]\times[-L/2,L/2]$ where $L$ = 100 nm illustrated in Fig. 2; periodic boundary conditions are applied throughout the boundary of the domain, while a difference of temperature of 1K is imposed along the x-axis. We model nano-scale heat transport via the Boltzmann transport equation under the relaxation time approximation, which, in the pseudo-temperature formulation reads [romano2021efficient]
$v_{\mu}\cdot\nabla T_{\mu}(r)=\dfrac{T_{\mu}(r)-T_{\mu}^{0}(r)}{\tau_{\mu}}\quad\quad\quad\quad\quad\quad\quad r\in\Omega,$ (2)
where $\mu$ collectively describes phonon wave vector and polarization. In Eq. 2, $\mathbf{v}_{\mu}$ and $\tau_{\mu}$ are the mode-resolved group velocity and scattering time, both computed by first-principled calculations and obtained by the AlmaBTE’s database [carrete2017almabte]. The term $T_{\mu}^{0}$ is an effective temperature, which is obtained by ensuring energy conservation; its expression is given by [romano2021efficient]
$f_{\mu}^{0}(r)=\sum_{\mu^{\prime}}\alpha_{\mu^{\prime}}f_{\mu^{\prime}}(r)$ (3)
@@
$\alpha_{\mu^{\prime}}=\frac{(C/\tau)_{\mu^{\prime}}}{\sum_{\mu^{*}}(C/\tau)_{\mu^{*}}}.$ (4)
@@
The bulk thermal conductivity is given by
$\kappa^{\alpha\beta}=\sum_{\mu}C_{\mu}^{\alpha}v_{\mu}^{\alpha}v_{\mu}^{\beta}\tau_{\mu},$ (5)
where $C_{\mu}$ is the mode-resolved heat capacity. In this work, we choose Si as the underlying material, with a thermal conductivity of $\kappa_{\text{bulk}}=150$W m-1 K-1. We parametrize the nanostructured material by a binary vector $\mathbf{p}$ of size 25, corresponding to a grid of 5x5 embedded pores of size 10x10 (nm) (example in Fig. 3B), yielding a total of approximately $33$ million ($2^{25}$) possible configurations. The pore walls diffusely scatter incoming phonons isotropically, leading to the boundary condition
$T_{B}=\frac{\sum_{\mu^{+}}C_{\mu}T_{\mu}\mathbf{v}\cdot\mathbf{\hat{n}}}{\sum_{\mu^{\prime}}C_{\mu^{+}}\mathbf{v}_{\mu}\cdot\mathbf{\hat{n}}}$, (6)
where $\mathbf{\hat{n}}$ is outward normal to the wall and $\mu^{+}$ refers to phonons incoming to the surface. Once Eq. 2 is solved, the effective thermal conductivity is computed by Fourier’s law
$\kappa=-\frac{1}{\Delta T}\int_{-L/2}^{L/2}\mathbf{J}(L/2,y)\cdot\mathbf{\hat{x}}dy,$ (7)
@@
where $\mathbf{J}=\sum_{\mu}C_{\mu}\mathbf{v}_{\mu}T_{\mu}$ is the heat flux. The goal of this work is to find a surrogate for the relationship $\kappa(\mathbf{p})$. With our surrogates we aim to achieve a 5.005.00% fractional error (FE) defined as $FE=\frac{|\kappa-\kappa_{\theta}|}{|\kappa|}$ with respect to the ground-truth effective thermal conductivity. Our ground truth is the conductivity obtained using the open-source software OpenBTE [openbte]. The low-fidelity solver used in this study is the Fourier solver developed by the OpenBTE team (yet to be released). To enhance computational efficiency, a coarser spatial conductivity discretization is adopted: a $5\times 5$ grid compared to the $100\times 100$ used in the BTE solver. This resolution is the minimum required to adequately represent the $5\times 5$ pore configuration.
@@ 4.1 Predictive Performance and Data Efficiency
We compared a vanilla MLP, PEDS and 4 models ensembles of these two architectures on a training dataset of randomly sampled geometry-conductivity pairs. The validation set and test set are composed of 1000 pairs, also randomly sampled. We trained for 1000 epochs with an ADAM optimizer and a cosine schedule with maximum learning rate of $5\cdot 10^{-3}$ and a minimum of $5\cdot 10^{-4}$. PEDS architecture was chosen to be relatively small with 2 fully-connected hidden layers of size 64 and a resolution of 1 (5x5 grid). Consistently with previous work [Pestourie_Mroueh_Rackauckas_Das_Johnson], smaller resolutions seemed not to yield any improvement in terms of accuracy. All activation functions are ReLU, except for the final hard-tanh layer, which constrains the output to the range $0,\kappa_{\text{bulk}}]$. All model parameters were initialized using Xavier-normal initialization. The learnable parameter $w_{\theta}(x)$ controlling the geometries linear combination is the output of a fully connected perceptron with a sigmoid activation and Kaiming initialization [he2015delving]. The MLP baseline architecture was chosen to resemble PEDS’, having an additional layer to match the number of parameters. Other number of layers, activation functions and initializations were tested before converging to these architectural choices. The training was parallelized and performed on 4 CPU cores. For completeness, we also included the performance of a Gaussian Process (GP) in our comparison. GPs are a widely used Bayesian approach for regression tasks [rasmussen2006gaussian]. Their flexibility and strong performance in low-data regimes make them a standard benchmark in surrogate modeling and scientific machine learning. We experimented with both RBF and Matern kernels, classic choices for modeling smooth and moderately rough functions respectively [rasmussen2006gaussian]. A Matern kernel with learned length scale and smoothness parameter, combined with epistemic white noise with inferred intensity yielded the best predictive accuracy.
@@
The ensemble of PEDS combined with active learning reduces the fractional test error by approximately 70% compared to PEDS without active learning and by about 75% compared to a purely data-driven MLP trained with active learning. A single PEDS model alone lowers the error by roughly 50% relative to a single MLP. These gains are most pronounced in the small-data regime, where PEDS starts well below data-driven baselines and reaches the 5% error threshold with substantially fewer training samples. Across random seeds, the GP serves as a stable, low-variance surrogate and performs surprisingly well as a general-purpose baseline, though its improvement from active learning is modest compared to PEDS. Mechanistically, the strong PEDS + AL performance arises from its model structure and sampling strategy: PEDS incorporates a low-cost, physics-based Fourier core to capture the governing physics, while the neural generator learns pore-scale corrections that achieves the surrogate’s accuracy. This reduces the complexity of the learning task and, consequently, the number of expensive BTE labels required for good generalization. AL amplifies those savings by directing costly simulations to the most informative data points. Empirically, we find the AL loop oversamples geometries in the tails of the effective-conductivity distribution. The model first captures the dominant posterior modes and then improves by incorporating these rarer, small-$\kappa$ and high-$\kappa$ examples. In contrast, a plain MLP must learn both macro- and micro-scale behavior from data and therefore requires many more examples, and GP, while sample-efficient and uncertainty-aware, cannot exploit the physics-informed structure to the same extent. Overall these results indicate that AL delivers the biggest marginal gains when it is applied to input-space generation and paired with a physics-informed PEDS model.This effect is consistent with classical active learning surveys and with surrogate-assisted optimization practice (e.g. EGO and other related strategies) [settles2012active, forrester2007multi].
@@
Figure 3: Panel (a) quantifies data efficiency by plotting fractional test error against training-set size for three surrogate approaches: PEDS-ENSEMBLE (blue), a MLP-ENSEMBLE (orange) and a Gaussian Process (green), showing results both with Active Learning (solid lines) and without (dash–dot). Error bars reflect standard deviations for the 5 seeds. The main result is that physics-enhanced learning with Active Learning produces the largest reductions: PEDS paired with AL lowers the fractional error by $\sim$ 70% relative to PEDS without AL, $\sim$75% relative to MLP with AL, and PEDS alone already reduces error by $\sim$ 50% compared with the MLP. The GP delivers steady and robust performance with small variance, but its improvement from AL is modest compared with PEDS. Panel (b) shows an optimized periodic pore pattern found by coupling a PEDS-ENSEMBLE surrogate and AL inside a Bayesian optimization loop to hit an effective conductivity of 30.00, illustrating how the combined framework discovers nontrivial, high-performing topologies with far fewer expensive OpenBTE evaluations. Means and standard deviations are computed on 5 repetitions.
Training
Evaluations
PEDS-ENS (%) PEDS+AL (%) GP (%) GP+AL (%) MLP-ENS (%) MLP+AL (%)

### v2

@@ 2.1 Uncertainty Quantification with Deep Ensembles and Heteroskedastic Variance
Uncertainty Quantification (UQ) is essential in surrogate modeling for scientific applications [73], especially when high-fidelity simulations are computationally expensive and data is limited [1]. We adopt a heteroskedastic Gaussian surrogate model, where the mean $\kappa_{\theta}(x)\in\mathbb{R}^{N}$ is predicted by an ensemble of PEDS [49] and the input-dependent log-variance $(\log\sigma^{2})_{\theta}(x)$ is predicted by a vanilla neural network. The model is trained using a Gaussian negative log-likelihood loss [7, 49]: $\mathcal{L}_{\text{NLL}}=\sum_{i}\beta\cdot(\log\sigma^{2})_{\theta}(x_{i})+(1-\beta)\cdot\frac{(\kappa(x_{i})-\kappa_{\theta}(x_{i}))^{2}}{\exp((\log\sigma^{2})_{\theta}(x_{i}))}$, where the parameter $\beta\in[0,1]$ controls the trade-off between penalizing overconfident predictions and fitting the data. We train an ensemble of $M$ independently initialized surrogate models to further capture uncertainty arising from model variability, especially from the optimization of their parametric parts. The ensemble prediction is given by the empirical mean $\hat{\kappa}(x)=\frac{1}{M}\sum_{i=1}^{M}\kappa_{i}(x)$, and the total predictive variance combines the predicted variance and the variance of the predictions as $\hat{\sigma}^{2}(x)=\frac{1}{M}\sum_{i=1}^{M}\sigma_{i}^{2}(x)+\frac{1}{M}\sum_{i=1}^{M}(\kappa_{i}(x)-\hat{\kappa}(x))^{2}$. This approach robustly predicts epistemic uncertainty, supports active learning by prioritizing uncertain samples, and facilitates stochastic optimization by taking into account the confidence in the surrogate’s output rather than relying solely on point estimates.
@@ 2.2 Active Learning
PEDS has previously shown great data savings and better performance on challenging designs when paired with an Active Learning framework [71]. We implemented an uncertainty-driven active learning pipeline using our uncertainty measure following Ref. [70]. We initialize a dataset with N training points, we train our surrogate and then propose M new geometries, compute their uncertainty and only compute and add to the dataset the $K\ll M$ most uncertain points. This is in contrast with other sample proposal strategies that are based upon the diversity of the samples or a mixture of diversity and uncertainty [81, 83]. In our case the measure of uncertainty is the variance as computed above. For active learning, unlike complex data assimilation problems requiring detailed distribution estimates, a coarse approximation of uncertainty is sufficient [1].
@@ 3 Design of Porous-structure with desired effective Thermal Conductivity
At the nano-scale, heat conduction diverges from the classical Fourier Law [15, 16] and assumes the form of a ballistic process rather than a diffusive one. This is because the mean free path (MFP) of heat-carrying quasi-particles in semiconductors, i.e. phonons, is comparable to the characteristic dimensions of the material. This phenomenon, often referred to as phonon-size effect [15, 89], leads, in porous material, to a suppressed effective thermal conductivity compared to the one obtained with a standard Fourier solver [87, 32, 50] (Fig. 2). These effects hold great promise across multiple engineering domains, shifting the focus from bulk materials to nanostructures and their inverse design.
In recent work, direct topology‐optimization of porous nanostructures has been attempted using the phonon BTE. For example, Ref. [94] employed a genetic‐algorithm search with a gray (single‐MFP) BTE model for nanoporous graphene and found disordered pore patterns that increase thermal conductivity. However, gray (single mean-free-path) models neglect the full phonon spectrum and generally overestimate conductivity. In fact, gray-model solutions can qualitatively differ from full-spectrum solutions: for instance, Ref. [36] shows that a gray BTE systematically predicts a higher effective conductivity than a mode-resolved BTE for the same nanofilm. Thus, designs based on single‐MFP models may yield very different results than those from a full phonon-spectrum BTE. The development of differentiable BTE frameworks is an active research area. The first implementation of such a framework in the context of inverse design (Ref. [77]) focused on the single-MFP model; only recently, a mode-resolved differentiable BTE solver (based on the relaxation-time approximation) was developed [82]; however, this framework did not include the interpolation model needed for inverse design (Ref. [77]).
@@ 3.1 Problem Statement
We consider a 2D domain $\Omega=[-L/2,L/2]\times[-L/2,L/2]$ where $L$ = 100 nm illustrated in Fig. 2. We model nano-scale heat transport via the Boltzmann transport equation under the relaxation time approximation, which in the pseudo-temperature formulation [78], reads
$v_{\mu}\cdot\nabla T_{\mu}(\mathbf{r})=\dfrac{T_{\mu}(\mathbf{r})-\bar{T}(\mathbf{r})}{\tau_{\mu}}\quad\quad\quad\quad\quad\quad\quad\mathbf{r}\in\Omega,$ (2)
where $T_{\mu}(\mathbf{r})$ is the phonon distribution for mode $\mu$, which collectively describes wave vector and polarization. The term $\mathbf{v}_{\mu}$ is the mode-resolved group velocity, and $\tau_{\mu}$ denotes the scattering time for mode $\mu$, which includes three-phonon as well as naturally occurring isotope scattering [98]. The term $\bar{T}(\mathbf{r})$ is the pseudo-temperature [29], which is the quantity plotted in the examples. Its expression is [79]
$\bar{T}(\mathbf{r})=\sum_{\mu^{\prime}}\alpha_{\mu^{\prime}}T_{\mu^{\prime}}(\mathbf{r})$ (3)
@@
$\alpha_{\mu^{\prime}}=\frac{(C/\tau)_{\mu^{\prime}}}{\sum_{\mu^{{}^{\prime\prime}}}(C/\tau)_{\mu^{{}^{\prime\prime}}}}.$ (4)
In Equation 4, $C_{\mu}$ is the mode-resolved volumetric heat capacity, which also include the normalization factor arising from real-space discretization. We note that both $T_{\mu}(\mathbf{r})$ and $\bar{T}(\mathbf{r})$ are deviations from a reference temperature $T_{0}$, which we conveniently set to 0 K. Throughout this work, we refer to $\bar{T}(\mathbf{r})$ as simply the temperature. Within this formulation, the flux is
$\mathbf{J}(\mathbf{r})=\sum_{\mu}C_{\mu}\mathbf{v}_{\mu}T_{\mu}(\mathbf{r}).$ (5)
It is straightforward to show that after including Eq. 3 into Eq. 2 and using Eq. 5, we obtain $\nabla\cdot\mathbf{J}(\mathbf{r})=0$, which is the usual steady-state continuity equation. The pore walls, denoted here by $\mathbf{r}_{B}$, diffusely scatter incoming phonons isotropically (hard-wall boundary condition), such that outgoing phonon modes $\mu^{\prime}$ (i.e., those with $\mathbf{v}_{\mu^{\prime}}\cdot\mathbf{\hat{n}}>0$, where $\mathbf{\hat{n}}$ is the outward normal to the wall) satisfy
$T_{\mu^{\prime}}(\mathbf{r}_{B})=\frac{\sum_{\mu^{+}}C_{\mu}T_{\mu}(\mathbf{r}_{B})\mathbf{v}_{\mu^{+}}\cdot\mathbf{\hat{n}}}{\sum_{\mu^{+}}C_{\mu^{+}}\mathbf{v}_{\mu^{+}}\cdot\mathbf{\hat{n}}},\quad\mathbf{v}_{\mu^{\prime}}\cdot\mathbf{\hat{n}}>0,$ (6)
where $\mu^{+}$ refers to phonon modes incoming to the surface (i.e., $\mathbf{v}_{\mu^{+}}\cdot\mathbf{\hat{n}}<0$) and $\mu^{\prime}$ denotes the reflected (outgoing) phonon directions. Periodic boundary conditions are applied throughout the simulation domain, where a jump of temperature jump $\Delta T=1K$ is also applied across the $y$-axis. The resulting boundary conditions are
$T_{\mu}\left(-L/2,y\right)$ $=$ $T_{\mu}\left(L/2,y\right),$
$T_{\mu}\left(x,L/2\right)$ $=$ $T_{\mu}\left(x,-L/2\right)+\Delta T,$ (7)
Since $\Delta T$ is isotropic and $\sum_{\mu}C_{\mu}\mathbf{v}_{\mu}=0$ is zero due to time-reversal symmetry, Eq. 3.1 ensures the continuity of the heat flux across the unit-cell border. Furthermore, combining Eqs. 3-4 with Eq. 3.1, we have $\bar{T}(-L/2,y)=\bar{T}(L/2,y)$ and $\bar{T}(x,L/2)=\bar{T}(x,-L/2)+\Delta T$, where we used $\sum_{\mu^{\prime+(-)}}\alpha_{\mu^{\prime}}=1/2$. Therefore, the imposed $\Delta T$ at the phonon-mode level translates into a jump in $\bar{T}(\mathbf{r})$. We note that this form of periodic boundary conditions are commonly used in BTE solvers (e.g., see [34])
Once Eq. 2 is solved, the effective thermal conductivity, which is our ground truth, is computed by Fourier’s law
$\kappa=-\frac{1}{\Delta T}\int_{-L/2}^{L/2}\mathbf{J}(x,L/2)\cdot\mathbf{\hat{y}}dx,$ (8)
@@
In absence of boundaries, the effective thermal conductivity equals the bulk one, given by
$\kappa^{\alpha\beta}=\sum_{\mu}C_{\mu}v_{\mu}^{\alpha}v_{\mu}^{\beta}\tau_{\mu},$ (9)
which is the standard kinetic formula [98]. In this work, we choose Si as the underlying material, which gives $\kappa_{\text{bulk}}\approx 150$W m-1 K-1. The second- and third-order force constants are obtained from the AlmaBTE database [14], which computes the forces based on the supercell approach and density functional theory [9]. The scattering times are computed on a reciprocal-space grid of $32\times 32\times 32$. Equation 2 is solved using the open-source software OpenBTE [80], which discretizes the spatial domain with the finite-volume technique, and bins the reciprocal space into a set of directional MFPs [79]. In this work, we employ 50 MFPs and 48 polar angles. Details about the solver implementation can be found in [80]
@@
We parametrize the nanostructured material by a binary vector $\mathbf{x}$ of size 25, corresponding to a grid of 5x5 embedded pores of size 10x10 (nm) (example in Fig. 3B), yielding a total of approximately $33$ million ($2^{25}$) possible configurations. The goal of this work is to find a surrogate for the relationship $\kappa(\mathbf{x})$. With our surrogates we aim to achieve a 5.005.00% fractional error (FE) defined as $FE=\frac{|\kappa-\kappa_{\theta}|}{|\kappa|}$ with respect to the ground-truth effective thermal conductivity. The low-fidelity solver used in this study is the Fourier solver developed by the OpenBTE team (yet to be released). To enhance computational efficiency, a coarser spatial conductivity discretization is adopted: a $5\times 5$ grid compared to the $100\times 100$ used in the BTE solver. This resolution is the minimum required to adequately represent the $5\times 5$ pore parametrization. The Fourier solver takes as input a spatially varying thermal conductivity field, where the thermal conductivity within pore regions is set to zero. The governing equation reads
$\nabla\cdot\left[\kappa(\mathbf{r})\nabla T(\mathbf{r})\right]=0,\quad\mathbf{r}\in\Omega,$ (10)
with adiabatic boundary conditions
$\kappa\nabla T(\mathbf{r})\cdot\mathbf{\hat{n}}=0\quad\quad\quad\quad\mathbf{r}\in\mathbf{r}_{B},$ (11)
and periodic boundary conditions
$T(-L/2,y)$ $=$ $T(L/2,y)$
$T(x,L/2)$ $=$ $T(x,-L/2)+\Delta T.$ (12)
Similarly to the BTE case, the flux $\kappa\nabla T(\mathbf{r})$ is continuous across the border because $\nabla\Delta T=\mathbf{0}$.
@@ 4.1 Predictive Performance and Data Efficiency
We compared a vanilla MLP, PEDS, and 4-model ensembles of both architectures. The training dataset was generated by randomly sampling binary 25-dimensional parametrizations and running the openbte mode-resolved steady-state solver on the resulting geometries; In our base experiment the training dataset was chosen to be 1000 pairs. Similarly, the validation and test sets each consist of 1000 randomly sampled pairs. No preprocessing was applied to the inputs or outputs beyond what is described below. We trained for 1000 epochs with an ADAM optimizer and a cosine schedule with maximum learning rate of $5\cdot 10^{-3}$ and a minimum of $5\cdot 10^{-4}$. PEDS architecture was chosen to be relatively small with 2 fully-connected hidden layers of size 64 and a resolution of 1 (5x5 grid). Consistently with previous work [71], smaller resolutions seemed not to yield any improvement in terms of accuracy. All activation functions are ReLU, except for the final hard-tanh layer, which constrains the output to the range $0,\kappa_{\text{bulk}}]$. All model parameters were initialized using Xavier-normal initialization. The learnable parameter $w_{\theta}(x)$ controlling the geometries linear combination is the output of a fully connected perceptron with a sigmoid activation and Kaiming initialization [31]. The MLP baseline architecture was chosen to resemble PEDS’, having an additional layer to match the number of parameters. Other number of layers, activation functions and initializations were tested before converging to these architectural choices. The training was parallelized and performed on 4 CPU cores. For completeness, we also included the performance of a Gaussian Process (GP) in our comparison. GPs are a widely used Bayesian approach for regression tasks [76]. Their flexibility and strong performance in low-data regimes make them a standard benchmark in surrogate modeling and scientific machine learning. We experimented with both RBF and Matern kernels, classic choices for modeling smooth and moderately rough functions respectively [76]. A Matern kernel with learned length scale and smoothness parameter, combined with epistemic white noise with inferred intensity yielded the best predictive accuracy.
@@
The ensemble of PEDS combined with active learning reduces the fractional test error by approximately 70% compared to PEDS without active learning and by about 75% compared to a purely data-driven MLP trained with active learning. A single PEDS model alone lowers the error by roughly 50% relative to a single MLP. These gains are most pronounced in the small-data regime, where PEDS starts well below data-driven baselines and reaches the 5% error threshold with substantially fewer training samples. Across random seeds, the GP serves as a stable, low-variance surrogate and performs surprisingly well as a general-purpose baseline, though its improvement from active learning is modest compared to PEDS. Mechanistically, the strong PEDS + AL performance arises from its model structure and sampling strategy: PEDS incorporates a low-cost, physics-based Fourier core to capture the governing physics, while the neural generator learns pore-scale corrections that achieves the surrogate’s accuracy. This reduces the complexity of the learning task and, consequently, the number of expensive BTE labels required for good generalization. AL amplifies those savings by directing costly simulations to the most informative data points. Empirically, we find the AL loop oversamples geometries in the tails of the effective-conductivity distribution. The model first captures the dominant posterior modes and then improves by incorporating these rarer, small-$\kappa$ and high-$\kappa$ examples. In contrast, a plain MLP must learn both macro- and micro-scale behavior from data and therefore requires many more examples, and GP, while sample-efficient and uncertainty-aware, cannot exploit the physics-informed structure to the same extent. Overall these results indicate that AL delivers the biggest marginal gains when it is applied to input-space generation and paired with a physics-informed PEDS model.This effect is consistent with classical active learning surveys and with surrogate-assisted optimization practice (e.g. EGO and other related strategies) [81, 22]. The data efficiency of PEDS is unprecedented compared to prior hybrid physics-ML surrogates  [58]. In fact, to compute the thermal conductivity, also accounting for its multi-fidelity version, the DeepONet strategy in [58] required a significantly larger number of evaluations (10x with respect to our base experiment, almost 30x if we consider our active learning). Nonetheless, these two strategies are not perfectly comparable. In principle DeepONet is a solver, and this reflects also on the choice of computing the error on the temperature field rather than on the scalar effective conductivity.
@@
Figure 3: Panel (a) quantifies data efficiency by plotting fractional test error against training-set size for three surrogate approaches: PEDS-ENSEMBLE (blue), a MLP-ENSEMBLE (orange) and a Matern kernel GP (green), showing results both with Active Learning (solid lines) and without (dash–dot). Error bars reflect standard deviations for the 5 seeds. The main result is that physics-enhanced learning with Active Learning produces the largest reductions: PEDS paired with AL lowers the fractional error by $\sim$ 70% relative to PEDS without AL, $\sim$75% relative to MLP with AL, and PEDS alone already reduces error by $\sim$ 50% compared with the MLP. The GP delivers steady and robust performance with small variance, but its improvement from AL is modest compared with PEDS. Panel (b) shows an optimized periodic pore pattern found by coupling a PEDS-ENSEMBLE surrogate and AL inside a Bayesian optimization loop to hit an effective conductivity of 30.00.
Training Evaluations PEDS-ENS (%) PEDS+AL (%) GP (%) GP+AL (%) MLP-ENS (%) MLP+AL (%)